OpenStudy (anonymous):

a man wishes to fence in a rectangular enclosure of area 128m^2.one side of the enclosure is formed by part of the brickwall already in place .what is the least possible length of fencing required for the other three sides?

OpenStudy (anonymous):

answer to a square is 33.941125496954281171240529381033 but if one side is 16 then you only need 32 of fencing

myininaya (myininaya):

A=xy=128 xy=128 y=128/x assuming y is the length brickwall then the side parallel to has same length P=y+2x P=128/x+2x P'=-128/(x^2)+2 set P'=0 -128/(x^2)=-2 128/(x^2)=2 128=2x^2 64=x^2 x=8 y=128/8=16

OpenStudy (anonymous):

myininaya what happen to your solution? one side that you had as 128/sqrt(128) is what I got. and I miss reading the problem? that works if a square but i tried a random side of 16 and found 128/16 = 8 therefore only 32 feet needed

myininaya (myininaya):

we dont want fencing on the brickwall

myininaya (myininaya):

we want fencing on 3 sides not 4

OpenStudy (anonymous):

I agree but is the question the total length of fence or keeping each side the shortest?

myininaya (myininaya):

so the answer is P=y+2x=16+2(8)=16+16=32

myininaya (myininaya):

it is the sum of the 3 side's lengths when we have found the shortest lengths possible

OpenStudy (anonymous):

I agree with your answer but I missed something with the question. I got the answer just because I had seen enough of these problems and knew to factor 128 in to common numbers 2,64 4,32 8,16 and so on.

myininaya (myininaya):

i assume you taken calculus

myininaya (myininaya):

we want to minimize

myininaya (myininaya):

so we need P'

OpenStudy (anonymous):

thanx guys

myininaya (myininaya):

and set =0

myininaya (myininaya):

to find critical numbers

myininaya (myininaya):

there is only one critical number and it is a minimum

OpenStudy (anonymous):

true but why use a sledgehammer to kill a fly? most are looking for answers not the correct method. that is why I checked your answer as the best because the solution was given.

myininaya (myininaya):

lol

myininaya (myininaya):

dont you think it is funner to kill a fly with a sledgehammer